### v1

# Shotluck Holmes: A Family of Efficient Small-Scale Large Language Vision Models For Video Captioning and Summarization

Abstract
Video is an increasingly prominent and information-dense medium, yet it poses substantial challenges for language models. A typical video consists of a sequence of shorter segments, or shots, that collectively form a coherent narrative. Each shot is analogous to a word in a sentence where multiple data streams of information (such as visual and auditory data) must be processed simultaneously. Comprehension of the entire video requires not only understanding the visual-audio information of each shot but also requires that the model links the ideas between each shot to generate a larger, all-encompassing story. Despite significant progress in the field, current works often overlook videos’ more granular shot-by-shot semantic information. In this project, we propose a family of efficient large language vision models (LLVMs) to boost video summarization and captioning called Shotluck Holmes. By leveraging better pretraining and data collection strategies, we extend the abilities of existing small LLVMs from being able to understand a picture to being able to understand a sequence of frames. Specifically, we show that Shotluck Holmes achieves better performance than state-of-the-art results on the Shot2Story video captioning and summary task with significantly smaller and more computationally efficient models.
Figure 1: A qualitative demo of the 3.1B model of Shotluck Holmes. Here, we demonstrate that our model is able to summarize data from many different parts of a long video using both visual and audio data.
1 Introduction
Over the last five years, the capability for machine learning models to intake, understand, and critically reason with natural language and visual data has exploded, primarily due to advances in model architecture Vaswani et al. (2017), compute capability, and a huge increase in available data. In particular, multi-modal large language models (LLMs), powered by the revolutionary transformer architecture Vaswani et al. (2017) have been able to achieve record-breaking understanding and reasoning capabilities on natural language and audiovisual understanding. Due to their resounding success in these fields OpenAI (2024), LLMs have also been at the forefront in the effort for AI models to understand video. Video is incredibly complex, since it combines dynamic movement in a visual medium with aural narration, sounds, text. These four aspects tend to constructively and destructively interfere with each other over the course of a given video. As such, it remains challenging for state-of-the-art (SOTA) language models to effectively comprehend and reason off of them. The current SOTA approach, Shot2Story20K Han et al. (2023), proposes a landmark new benchmark dataset that combines visual and auditory signals through a three-stage model pipeline. Then, they use this custom benchmark dataset to train a custom model architecture. This approach greatly improved performance on single-shot narration captioning, multi-shot video summarization, video Q&\&&A, and video retrieval, showing that embedding multiple sources of information is essential for language models to gain an improved understanding of these multi-modal inputs.
In this paper, we take the advances presented by the Shot2Story20K paper and integrate it with one of the leading small-scale multi-modal model families: TinyLLaVA. Specifically, we show that it is sufficient to replace the final two stages of Shot2Story20K’s three-stage vision-language model pipeline with TinyLLaVA, and that this replacement not only greatly reduces the memory usage, compute footprint, and latency, but also achieves SOTA performance (despite the much smaller model size) after finetuning on the Shot2Story20K dataset.
2 Relevant Work
Large Multimodal Models.    Due to their impressive capabilities, large language models (LLMs) have garnered significant research interest in recent years Chowdhery et al. (2023); Chiang et al. ; Gunasekar et al. (2023); Zhang et al. (2024); Brown et al. (2020). This, combined with advancements in vision encoders Zhai et al. (2023); Sun et al. (2023), has led to some significant works in the multimodal Large Language Model field. Recent LLMs such as LLaVA Liu et al. (2024) and InstructBLIP Dai et al. (2024) leverage fine-tuning on existing LLM backbones with visual instruction tuning data to improve zero-shot performance and model alignment with human preferences. However, these models are rather large at 7B and 8.2B parameters respectively.
TinyLLaVA.    The TinyLLaVA framework provides analysis on exploiting various small-scale LLMs for LLVMs, utilizing the same fine-tuning approach on visual instruction tuning data. Their research Zhou et al. (2024) finds that SigLIP Zhai et al. (2023) yields better performance than CLIP when combined with small-scale LLMs of varying parameter ranges, including TinyLlama (1.1B) Zhang et al. (2024) and Phi-2 (2.7B) Li et al. (2023b). It was found that small-scale LLVMs (3.1B) can achieve better overall performance against existing 7B models including LLavA-1.5 and Qwen-VL on various evaluation benchmarks such as TextVQA Singh et al. (2019), SQA-I Lu et al. (2022), GQA Hudson and Manning (2019), VQAv2 Goyal et al. (2017), MMB Liu et al. (2023), MME Yin et al. (2023), LLaVA-W Liu et al. (2024), POPE Li et al. (2023a), and M-Vet Yu et al. (2023). However, the capabilities of TinyLLaVA are limited to just image-text-to-text generation.
Shot2Story20K.    Shot2Story20K Han et al. (2023) leverages HDvila100M Xue et al. (2022) (which contains abundant automatic speech recognition (ASR) content) and performs strict quality checking to collect a rich dataset of size 20023. A comparison of Shot2Story20K with pre-existing datasets (see Table 1) shows that Shot2Story20K outperforms its predecessors in quality. Recently, the authors have even released a 134K version of the dataset, placing this benchmark shot-level dataset far beyond the capabilities of its competitors.
Table 1: A high-level comparison of Shot2Story20K with pre-existing video-based datasets Han et al. (2023). M and G stand for manual and generated, respectively.
Dataset Annot.
Multi-shot
Video
Multi-event
Descriptions
Audio
Captions
Detailed
Summary
Summary
Length
# Vids.
Avg.
Duration
MSRVTT M - 10K 15s
ActivityNet Caps M 52.4 20K 3min
VideoStorytelling M 162.6 105 12.5min
Ego4D M - 10K 23min
YouCook2 M 67.8 2K 6min
VAST G 32.4 27M 5$\sim$30s
Shot2Story20K M+G 201.8 20K 16.7s
As mentioned, Shot2Story20K Han et al. (2023) combines visual and auditory signals by creating a dataset through a three-stage model. (1) TransNetV2 Souček and Lokoč (2020) separates videos into shots. (2) MiniGPT-4 Zhu et al. (2023) performs video captioning of individual shots and the results are manually checked by human annotators, who also write narration captions of individual shots. (3) GPT-4 OpenAI (2024) combines all individual shot captions and ASR to perform full video summarization with human prompting.
3 Method
3.1 Data-preprocessing
Multiple preprocessing steps were taken to adapt the video-based Shot2Story20K dataset for compatibility with TinyLLaVA. The Shotluck Holmes preprocessing procedure involved segmenting the Shot2Story20K annotations into entries comprising individual shots, each paired with its corresponding caption and ASR transcription. Subsequently, the annotations underwent processing to identify and eliminate corrupted data in order to remove any such corrupted shots from the dataset. The resultant Shotluck Holmes dataset is structured to include video inputs alongside a compilation of dialogues exchanged between a human prompt and the LLM ground truth output.
3.2 Video to Tensor
In order for the LLM to process the video data, the video is first converted into a tensor and then fed into a vision encoder such as SigLip Zhai et al. (2023). To convert the video into a tensor, we experimented with two sampling methods inspired by the approach in LAVIS Li et al. (2022): uniform sampling and head-tail sampling. Head-tail sampling forces the random sampling to sample an equal number of frames from the first half of the video and the second half of the video. Sampled frames are then concatenated and fed into a vision encoder.
3.3 Shotluck Holmes Backbones
The Shot2Story20K model architecture uses multiple different language models to assist in the video summarization or Q&\&&A tasks. By leveraging more efficient small-scale LLVMs like TinyLLaVA, we hope to achieve strong performance on single-shot video captioning and multi-shot video summarization with significantly reduced computational complexity by replacing the entire Shot2Story20K model architecture with a single LLVM model Zhou et al. (2024). Because TinyLLaVA models are already pre-trained with vision encoders Zhai et al. (2023) and small-scale language models, our goal is to utilize Shot2Story20K’s dataset and fine-tune the smaller LLM to extract similar performance as compared to Shot2Story’s multi-step model.
Figure 2: Shotluck Holmes model architecture Zhou et al. (2024)
Shotluck Holmes presents a family of small LLMs finetuned on the Shot2Story20K dataset. We present our first two models: a 1.5B parameter LLM and a 3.1B LLM both based on TinyLLaVA, as shown in Table 2.
Table 2: Size of baseline models
Model Size LLM Vision Encoder
Shot2Story 7B Vicuna BLIP
Shotluck-Holmes (1.5B) 1.5B TinyLlama SigLIP
Shotluck-Holmes (3.1B) 3.1B Phi-2 SigLIP
Our model architecture is shown in Figure 2. We follow TinyLLaVA’s original pipeline of feeding in visual data (which in our case requires additional processing of the video) into the vision encoder, which is then mapped by a MLP into the LLM embedding space. During our fine-tuning, we freeze the first 12 layers of the vision encoder and update the rest of the model.
4 Experiments
For both single-shot video captioning and video summarization, we follow the same instruction template as described in Appendix A. Note that fine-tuning in all scenarios is supervised and thus includes the ground truth, which is removed during evaluation. Our optimizer and finetuning hyperparameters are listed in Appendix B.
4.1 Single-Shot Video Captioning
The goal of single-shot video captioning is to generate descriptions for individual video shots (i.e. sections of a full video). The model first samples $N=120$ frames from a video shot using one of two sampling methods as described earlier. We chose $N=120$ based on the largest number of frames that we could feasiblty fit into our training hardware. These frames are then concatenated and fed into the vision encoder to produce visual tokens. The tokens are concatenated to a predetermined text prompt based on the type of LLM being used (see Appendix A), which is followed by the ASR text as additional context clues. Finally, this tensor is fed through the small-scale LLM of the given size to generate the caption for the video shot. Shot2Story20K discovered that applying ASR text as additional context cues in the text prompt results in slightly improved BLEU and METEOR metrics, with slight degradation in ROUGE. As a result, we follow this method and apply ASR text to our pipeline as well. The results are quite successful. For example, given the challenging shot below in Figure 3, the model generates the sentence: "In the video, a man in a black suit and tie is standing in front of a large screen displaying a boxing match. He is speaking into a microphone in front him."
Figure 3: The sampled frame for single-shot captioning
4.2 Multi-Shot Video Summarization
Multi-shot video summarization involves providing a rational summary describing a progression of events across different shots taken from the same video. We use a similar approach to that of single-shot captioning when training for multi-shot video summarization - our dataset included both single-shot and multi-shot video examples and the model was simultaneously finetuned to do both. Despite this lack of focus in finetuning, the model performs feasibly well as it is able to generate coherent summaries of multi-shot videos.
For example, for a video which contains different shots of a tablet computer and narration describing its specifications and features, the model generated the following summary: "The video is about a tablet computer that has a lot of features. It has a 64 gigabyte hard drive with two gigabytes of memory, a fast Intel Atom z3740 5F quad core data processor, and a detachable magnetic hinge that allows it to be easily connected to a keyboard. The tablet also has a latch list, which is a feature that allows you to quickly and easily engage and disengage the tablet from its keyboard".
4.3 Evaluation
Table 3: Performance of best models on single-shot video captioning
Model BLEU METEOR ROUGE CIDER
Shot2Story (7B+) 10.7 16.2 29.6 37.4
Shotluck-Holmes (3.1B) 8.7 25.7 36.2 63.2
Shotluck-Holmes (1.5B) 9.3 25.3 36.3 68.9
Figure 4: Single-shot video captioning results
We evaluate our model on the task of video-captioning using BLEU@4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), and ROUGE (Lin, 2004) on both the single-shot video captioning task and the multi-shot video summarization task. Our decoding method is top_p sampling (see Table 10 in Appendix B for numbers). We make sure to normalize our scores using the same approach as Shot2Story20K for comparability and consistency with the SOTA model’s metrics. We see that on the single-shot video captioning taks, the 1.5B model of Shotluck-Holmes’s 1.5B parameter model is competitive with the Shot2Story model, despite ours being around 78% smaller. Furthermore, other than the BLEU-4 metric, which evaluates by comparing the output text sample to the baseline text sample in a manner more conducive to evaluating translations, Shotluck-Holmes 1.5B exceeds the performance of Shot2Story by between 50 and 100%.
These performance gains are corroborated by the 3.1B model, which matches or improves on the gains seen by the 1.5B model.
Table 4: Performance of best models on multi-shot video summarization
Model BLEU METEOR ROUGE CIDER
Shot2Story (7B+) 11.7 19.7 26.8 8.6
Shotluck-Holmes (3.1B) 7.67 23.2 43 152.3
Shotluck-Holmes (1.5B) 6.48 21.3 40.2 144.3
Figure 5: Multi-shot video summarization results
We also evaluate the performance of Shot2Story vs. the two Shotluck-Holmes models on the multi-shot video summarization task using the same four metrics as mentioned above. The two models one again achieve comparable or superior performance to the larger Shot2Story model, though their improvements are slightly muted as compared to the single-shot video captioning task, achieving gains between 30 and 80% when compared to the baseline.
4.4 Qualitative Evaluation
Besides systematically evaluating on public benchmarks, we further qualitatively examined how well Shotluck Holmes performed on the video captioning tasks. As illustrated in Figure 1, Shotluck Holmes 3.1B was able to summarize information chronologically from the video. It was quite accurate at transcribing the narration and did not miss any key pieces of information. In Figure 3, the model was able to infer the context of the single-shot video (a boxing match) without explicitly being told that info in a narration.
4.5 Ablation Study: Sampling Methods
To understand the true underlying benefits of the new Shotluck-Holmes pipeline, we performed an ablation study to study the effects of different frame sampling strategies on model performance. Specifically, we aimed to investigate the impact of sampling methods on the performance of the model and the generalizability of our approach.
Table 5: Sampling Method Ablation Study on Shotluck-Holmes 1.5B using the Single-Shot Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (1.5B) w/ Uniform Sampling 9.3 25.3 36.3 68.9
Shotluck-Holmes (1.5B) w/ Head-Tail Sampling 8.2 25.5 36.5 65.0
Table 6: Sampling Method Ablation Study on Shotluck-Holmes 3.1B using the Single-Shot Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (3.1B) w/ Uniform Sampling 8.7 25.7 36.2 63.2
Shotluck-Holmes (3.1B) w/ Head-Tail Sampling 8.8 25.9 37.1 67.6
Table 7: Sampling Method Ablation Study on Shotluck-Holmes 1.5B using the Full Video Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (1.5B) w/ Uniform Sampling 6.48 21.3 40.2 144.3
Shotluck-Holmes (1.5B) w/ Head-Tail Sampling 6.40 21.4 41.0 151.9
Table 8: Sampling Method Ablation Study on Shotluck-Holmes 3.1B using the Full Video Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (3.1B) w/ Uniform Sampling 7.67 23.2 43 152.3
Shotluck-Holmes (3.1B) w/ Head-Tail Sampling 7.60 23.0 42.4 157.9
To investigate this, we studied two common video frame sampling techniques. The first technique, known as uniform sampling, gives each frame a equal probability to be selected. The dataloader then selected $N$ frames at random (in our experiments, we kept $N=120$, for consistency). In comparison, head-tail sampling disproportionately weights the likelihood of drawing frames from the beginning and end of a given shot or video to get a total of $N$ sampled frames. The results of each, evaluated using the standard four metrics, are in the four tables above.
As demonstrated by the studies, there is no significant trend observable between the two sampling techniques. In some models and tasks, head-tail sampling performed better, while in others, uniform sampling performed better. Nevertheless, the two sampling methods resulted not only in comparable results, but also both soundly beat the SOTA Shot2Story20K model, demonstrating that our approach is generalizable across sampling methods.
5 Conclusion
In this paper, we propose Shotluck-Holmes, a family of efficient models that achieve state-of-the-art performance on shot-level and full-length video understanding. This result is achieved by combining Shot2Story’s Han et al. (2023) multi-model pipeline, which integrates shot-level video annotations with audio-visual elements, with small-scale LLMs like TinyLLaVA. Furthermore, we demonstrate that these smaller models are able to achieve generalization capabilities on video captioning and summarization tasks that are competitive with larger models, suggesting that such a task is executable even on limited hardware or on edge devices.
Despite this, there is still room for improvement with regards to the training pipeline and computational resources. One major limitation is that we trained the model on a hybrid dataset of single-shot and multi-shot videos. Although this allows for the model to generalize well across both tasks with limited samples, it prevents us from specialized fine-tuning on each particular task, which could result in even greater performance gains.
In addition, due to compute accessibility restrictions while training and evaluating, we were forced to shard the dataset and distribute the training across multiple nodes, and it’s possible that the learning rate scheduler was not set up correctly for this sharded process.
Finally, to better represent the sequential nature of full-length video as a sequence of singular shots, a conversation feed over the shots would be required. This would allow for a proper mapping of attention over the sequence of shots. As a result, it’s possible that our approach may overemphasize a single or a series of shots in a multi-shot video.
6 Acknowledgements
We would like to thank Dr. Wei Xu and her teaching assistants (Anton Lavrouk, Marcus Ma, Andrew Li) for the course CS 7650: Natural Langauge Processing at Georgia Tech, from which this project was inspired. In addition, we would like to extend a major thank you to PACE @ Georgia Tech PACE (2017) for providing the plentitude of computational resources and GPUs needed to train our models.
Appendix A Instruction Templates
This is our instruction template for prompting. During supervised training, we include the ground truth in assistant, but for evaluation, this ground truth is removed.
{mdframed}
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: $<Video>$
Please describe this video. Do not include details that you are not sure of. This is what the speech in the video is saying: $<ASR>$ ASSISTANT: $<GroundTruth>$
Appendix B Hyperparameters
Table 9: Hyperparameters for Shotluck Holmes training
Optimizer Global Batch Size Learning Rate Epochs Max Length Weight Decay
Adam8bit 128 2e-5 1 3072 0
Table 10: Evaluation Parameter Settings
Parameter Value
temperature 0.2
top_p 0.9
no_repeat_ngram_size 3
Appendix C Compute Resources
All models were trained in an environment with 2 TB of RAM, 8x NVIDIA H100 GPUs, and 64 CPU cores. Fine tuning the 1.5B parameter model took approximately 6 hours and the 3.1B parameter model 8 hours with this setup.
Appendix D Broader Impact
Shotluck Holmes is based entirely on existing LLMs and datasets and introduces no new architectures or data. As such, our work inherits any existing limitations of LLMs and the Shot2Story20K dataset, including but not limited to hallucination and biased outputs. However, our work does not introduce any new implications for societal impact or require any new safeguards. Improved video summarization capabilities do not pose any greater risk of misuse than current technologies.